# Hybrid Memetic Search for Electric Vehicle Routing with Time Windows, Simultaneous Pickup-Delivery, and Partial Recharges

### v1

@@ VI Conclusion
This work investigated an emerging routing problem, namely EVRP-TW-SPD, and proposed a hybrid memetic algorithm called HMA to solve it. HMA integrates two novel components: PSSI for handling partial recharges and CDNS as the local search procedure. Both components can be easily applied to other EVRP variants. Comprehensive experiments showed that HMA outperformed existing methods across a wide range of instances. Moreover, a new, large-scale benchmark set was established from a real-world JD logistics application. This set contains 20 problem instances with 200, 400, 600, 800, and 1000 customers, each with 100 stations,
HMA can be further improved in several directions. First, the PSSI procedure could benefit from a learned function (e.g., as a neural network [35]) that evaluates insertions instead of heuristic rules, which might lead to higher-quality feasible routes within even smaller time budget. Second, dividing the original problem into smaller, manageable subproblems enhances efficiency and yields high-quality solutions. However, the success of this approach heavily depends on the effectiveness of the decomposition strategy. Rather than relying on hand-designed strategies, future improvements could leverage learning-based methods [36] or general-purpose pretrained models [37] to automatically design effective decomposition strategies tailored to specific customer distributions.
@@ Appendix A Proof of Proposition 1
Proof. Consider a feasible EVRP-TW-SPD solution $\bm{S}$. By definition, $\bm{S}$ satisfies all electricity, capacity, and time window constraints. Let $\bar{\bm{S}}$ be the solution obtained by removing all charging stations in $\bm{S}$. Suppose the travel time between any two nodes satisfies the triangle inequality, then for any two successive customers $c_{i}$ and $c_{j}$ in $\bar{\bm{S}}$ that are separated by a charging station $f$, the triangle inequality ensures:

### v2

@@ VI Conclusion
This work investigated an emerging routing problem, namely EVRP-TW-SPD, and proposed a hybrid memetic algorithm called HMA to solve it. HMA integrates two novel components: PSSI for handling partial recharges and CDNS as the local search procedure. Both components can be easily applied to other EVRP variants. Moreover, a new, large-scale benchmark set was established from a real-world JD logistics application. This set contains 20 problem instances with 200, 400, 600, 800, and 1000 customers, each with 100 stations. Comprehensive experiments showed that HMA outperformed existing methods across a wide range of instances.
Although HMA has shown excellent performance in the experiments, several promising directions exist for further enhancing HMA’s performance. First, the PSSI procedure could benefit from a learned function (e.g., as a neural network [37]) that evaluates insertions instead of heuristic rules, which might lead to higher-quality feasible routes within even smaller time budget. Second, for large-scale problems, dividing the original problem into smaller, manageable subproblems enhances efficiency and yields high-quality solutions. However, the success of this approach heavily depends on the effectiveness of the decomposition strategy. Rather than relying on hand-designed strategies, future improvements could leverage learning-based methods [38] or general-purpose pretrained models [39] to automatically design effective decomposition strategies tailored to specific customer distributions or constraint characteristics.
@@ Appendix A Proof of Proposition 1
Proof. Consider a feasible EVRP-TW-SPD solution $\bm{S}$. By definition, $\bm{S}$ satisfies all electricity, capacity, and time window constraints. Let $\bar{\bm{S}}$ be the solution obtained by removing all charging stations in $\bm{S}$. Suppose the travel time between any two nodes satisfies the triangle inequality; then, for any two successive customers $c_{i}$ and $c_{j}$ in $\bar{\bm{S}}$ that are separated by a charging station $f$, the triangle inequality ensures: